OpenStudy (anonymous):

4(7x - 1) = 3(2x + 1) - 4(x + 6)

OpenStudy (lieutenantgeneral):

26x + 17 = 0 x = -17/26

OpenStudy (justnick09):

\[28x-4=6x+1-4x+24\]

OpenStudy (justnick09):

\[28x-4=2x+25\]

OpenStudy (justnick09):

\[28x=2x+29\]

OpenStudy (justnick09):

\[26x=29\]